# Multilevel neural simulation-based inference

### v1

@@ 1 Introduction
Figure 1: Low- and high-fidelity cosmological simulations from the CAMELS data (Villaescusa-Navarro et al., 2023) studied in Section 5.4.
@@
Nevertheless, the initial training phase of neural SBI methods typically still requires a large number of simulations, preventing their application on more computationally expensive (and often more realistic) models which can take hours of compute time for simulating a single data-point. Examples include most tsunami models (Behrens and Dias, 2015), wind farm models (Kirby et al., 2023), nuclear fusion (Hoppe et al., 2021) and cosmology (Jeffrey et al., 2025) simulators. One potential approach to mitigating this issue is multi-fidelity methods (Peherstorfer et al., 2018): instead of relying on a single simulator, we often have access to a sequence of simulators with increasing computational cost and accuracy which we may be able to use to refine existing methods. This setting is quite common; for instance, for simulators requiring the numerical solution of ordinary, partial, or stochastic differential equations, where the choice of mesh size or step-size affects both the accuracy of the solution and the computational cost. Another example arises when modelling complex physical, chemical, or biological processes, where simplified approximations can be obtained by neglecting certain aspects of the system. This is exemplified in Figure 1, where a high-fidelity cosmological simulation includes baryonic astrophysics and thus appear smoother than the low-fidelity simulation.
@@
The key idea behind multi-fidelity methods is to leverage cheaper, less accurate simulations to supplement the more expensive, high-fidelity simulations, ultimately improving efficiency without sacrificing accuracy. This has been popular in the emulation literature since the seminal work of Kennedy and O’Hagan (2000), but applications to SBI are much more recent and include Jasra et al. (2019); Warne et al. (2018); Prescott and Baker (2020, 2021); Warne et al. (2022); Prescott et al. (2024), who proposed multi-fidelity versions of ABC. More recently, (Krouglova et al., 2025) also proposed a multi-fidelity method to enhance neural SBI approximations of the posterior based on transfer learning. However, their method is not supported by any theoretical guarantees, and our numerical experiments in Section 5.2 will demonstrate that it is sensitive to choices of hyperparameters and the paper does not provide any guidance for selecting these.
@@
In this paper, we propose a novel multi-fidelity method which is broadly applicable to neural SBI methods. Taking neural likelihood and neural posterior estimation as the main case-studies, we show that our approach is able to significantly reduce the computational cost of the initial training through the use of multilevel Monte Carlo (Giles, 2015; Jasra et al., 2020) estimates of the training objective. The approach has strong theoretical guarantees; under mild conditions, our main result (Theorem 1) directly links the reduction in computational cost to the accuracy of the low fidelity simulators, and we demonstrate (in Theorem 2) how to best balance the number of simulations at each fidelity level in the process. Importantly, and unlike Krouglova et al. (2025), our method has no additional hyperparameter. Our extensive experiments on models from finance, synthetic biology, and cosmology also demonstrate the significant computational advantages provided by our method.
@@ 2 Background
We first recall the basics of SBI methods and the related works on reducing computational cost in Section 2.1, then provide a brief introduction to multilevel Monte Carlo in Section 2.2.
@@ 2.1 Simulation-based inference
Let $\{\mathbb{P}_{\theta}:\theta\in\Theta\}$ be a parametric model family of distributions on some space $\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$ parameterised by $\theta\in\Theta\subseteq\mathbb{R}^{d_{\Theta}}$. We assume that the model is available in the form of a computer code, i.e. as a simulator, wherein simulating realisations $x\sim\mathbb{P}_{\theta}$ is straightforward, but the likelihood function $p(\cdot\,|\,\theta)$ associated with $\mathbb{P}_{\theta}$ is intractable. The generative process of simulators can be characterised by a pair $(\mathbb{U},G_{\theta})$, where $\mathbb{U}$ is a distribution (typically simple, such as a uniform or a Gaussian) on a space $\mathcal{U}\subseteq\mathbb{R}^{d_{\mathcal{U}}}$ which captures all of the randomness, and $G_{\theta}:\mathcal{U}\mapsto\mathcal{X}$ is a (deterministic) parametric map called the generator. Sampling $x\sim\mathbb{P}_{\theta}$ requires first sampling $u\sim\mathbb{U}$, and then applying the generator $x=G_{\theta}(u)$. In this paper, we consider the classical Bayesian inference problem of estimating the parameters $\theta$ of this simulator-based model given independent and identically distributed (iid) data $x_{1:m}^{o}=\{x_{j}^{o}\}_{j=1}^{m}\in\mathcal{X}^{m}$ collected from some true data-generating process. Specifically, we are interested in estimating the posterior distribution $\pi(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}p(x_{j}^{o}\,|\,\theta)\pi(\theta)$, where $\pi(\theta)$ is the prior distribution. As introduced below, this can for example be achieved via a neural SBI method which approximates the likelihood or posterior distribution.
@@
NLEs (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022; Radev et al., 2023a) are extensions of the synthetic likelihood approach (Wood, 2010; Price et al., 2018) that use flexible conditional density estimators, typically normalising flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), as surrogate models for the density of $\mathbb{P}_{\theta}$, i.e., the intractable likelihood function. The surrogate conditional density $q_{\phi}^{\text{NLE}}:\mathcal{X}\times\Theta\rightarrow[0,\infty)$ where $q_{\phi}^{\text{NLE}}(\cdot\,|\,\theta)$ is a density function for each $\theta\in\Theta$ and $\phi\in\Phi\subseteq\mathbb{R}^{d_{\Phi}}$ denotes its learnable parameters, is trained by minimising the negative log-likelihood with respect to $\phi$ on simulated samples. More precisely, let $\{(\theta_{i},x_{i})\}_{i=1}^{n}$ be the training data such that $\theta_{i}\sim\pi$ are realisations from the prior and $x_{i}\sim\mathbb{P}_{\theta_{i}}$ are realisations from the simulator, then $\hat{\phi}_{\text{MC}}:=\operatorname*{arg\,min}_{\phi\in\Phi}\ell_{\text{MC}}^{\text{NLE}}(\phi)$, where $\ell_{\text{MC}}^{\text{NLE}}$ is an empirical (Monte Carlo) estimate of negative expected log-density:
@@
Once the surrogate likelihood is trained, Markov chain Monte Carlo (MCMC) or variational inference methods are used to sample from the (approximate) posterior distribution $\pi_{\text{NLE}}(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}q^{\text{NLE}}_{\hat{\phi}_{\text{MC}}}(x^{o}_{j}\,|\,\theta)\pi(\theta)$. NLEs can therefore be regarded as being partially amortised—the surrogate $q_{\phi}^{\text{NLE}}$ need not be trained for a new observed dataset, however, MCMC needs to be carried out again. For a computationally costly simulator, we note that obtaining training samples can become a bottleneck, which affects the accuracy of estimating the expected loss. Thus, estimating the loss accurately with fewer samples is key to handling costly simulators.
@@
Instead of learning a surrogate likelihood, NPEs learn a mapping $x\mapsto p(\theta\,|\,x)$ from the data to the posterior using conditional density estimators that are based on normalising flow (Papamakarios and Murray, 2016; Radev et al., 2022), diffusion (Geffner et al., 2023; Sharrock et al., 2024; Gloeckler et al., 2024), or consistency models (Schmitt et al., 2024b). Similar to NLE, the conditional density $q_{\phi}^{\text{NPE}}:\Theta\times\mathcal{X}^{m}\rightarrow[0,\infty)$ is trained by minimising the negative log likelihood with respect to $\phi$ using data $\{(\theta_{i},x_{1:m,i})\}_{i=1}^{n}$ generated by first sampling from the prior $\theta_{i}\sim\pi$ and then the simulator $x_{1:m,i}=(x_{1,i},\ldots,x_{m,i})\sim\mathbb{P}_{\theta_{i}}$:
@@
Once $\phi$ is estimated, the NPE posterior is obtained as $\pi_{\text{NPE}}(\theta\,|\,x_{1:m}^{o})=q^{\text{NPE}}_{\hat{\phi}_{\text{MC}}}(\theta\,|\,x_{1:m}^{o})$. Although training $q_{\phi}^{\text{NPE}}$ incurs an upfront cost, this is a one-time cost as approximate posteriors for new observed datasets are obtained by a simple forward pass of $x_{1:m}^{o}$ through the trained networks, making NPEs fully amortised. Similar to NLE, the computationally costly step in NPE is the generation of training samples from running expensive simulators.
Note that both $q^{\text{NLE}}_{\phi}$ and $q^{\text{NPE}}_{\phi}$ usually include a summary function (often architecturally implicit in NPE). This is helpful when $\mathcal{X}$ is high-dimensional or the number of observations $m$ is large (Alsing et al., 2018; Radev et al., 2022), and for NPE it allows conditioning on datasets of different sizes.
Related work on reducing the cost of SBI methods.
We briefly note that beyond the aforementioned multi-fidelity methods, other works also aim to reduce the computational cost of SBI. In the context of ABC, adaptive sampling of the posterior using either sequential Monte Carlo techniques (Sisson et al., 2007; Beaumont et al., 2009; D. Moral et al., 2011) or Gaussian process surrogates (Gutmann and Corander, 2016; Meeds and Welling, 2014) has been a popular approach. A similar approach has been applied to neural SBI (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Papamakarios et al., 2019; Hermans et al., 2020; Durkan et al., 2020), where sequential training schemes are employed to reduce the number of calls to the simulator. Other works have tackled this problem using cost-aware sampling (Bharti et al., 2025), side-stepping high-dimensional estimation (Jeffrey and Wandelt, 2020), early stopping of simulations (Prangle, 2016), dependent simulations (Niu et al., 2023; Bharti et al., 2023), expert-in-the-loop methods (Bharti et al., 2022b), self-consistency properties (Schmitt et al., 2024a), parallelisation of computations (Kulkarni and Moritz, 2023), and the Markovian structure of certain simulators (Gloeckler et al., 2025). Our method, which employs multilevel Monte Carlo, can be combined with all of these compute-efficient SBI methods and is hence complementary to them.
@@ 3 Methodology
Recall that we are performing inference for a simulator $(G_{\theta},\mathbb{U})$ with a prior $\pi$ on the parameter $\theta\in\Theta$. We reparameterise the NLE and the NPE objective in terms of $u$ instead of $x$ (akin to the reparametrisation trick in variational inference), and express them more broadly using an arbitrary loss $\ell$ (to represent either $\ell^{\text{NLE}}$ or $\ell^{\text{NPE}}$) and an arbitrary function $f_{\phi}$ (to represent either $f^{\text{NLE}}_{\phi}$ or $f^{\text{NPE}}_{\phi}$) as:
@@
The MC estimator of the loss $\ell(\phi)$ is given by ${\ell}_{\text{MC}}(\phi):=\frac{1}{n}\sum_{i=1}^{n}f_{\phi}(u_{1:m,i},\theta_{i})$. Note that the variance of this MC estimator depends on the number of iid samples $n$. Hence, a small $n$ owing to a computationally expensive simulator will lead to a poor estimator for the loss. Now suppose that we have access to a sequence simulators (or generators) $G^{0}_{\theta}(u),\dots,G^{L-1}_{\theta}(u)$ with varying fidelity levels. Then, for $l=0,\ldots,L-1$, we can define a corresponding sequence of functions $f^{l}_{\phi}(u_{1:m},\theta)$:
$f^{\text{NLE},l}_{\phi}(u_{1:m},\theta):=-\log q^{\text{NLE}}_{\phi}\left(G^{l}_{\theta}(u)\,|\,\theta\right),\,\text{and}\,f^{\text{NPE},l}_{\phi}(u_{1:m},\theta):=-\log q^{\text{NPE}}_{\phi}\left(\theta\,|\,G^{l}_{\theta}(u_{1}),\ldots,G^{l}_{\theta}(u_{m})\right),$
@@
Equation 1 follows by adding then subtracting some terms, and using linearity of expectations. Now suppose that we can simulate from each of these approximate simulators to obtain:
@@
This corresponds to an MLMC estimator of our original objective in 1. The idea is that $h_{0}(\phi)$ is a cheap but rough estimate of $\ell_{\text{MC}}(\phi)$, and each additional term $h_{l}(\phi)$ accounts for the difference between the $l^{\textup{th}}$ and the $(l-1)^{\textup{th}}$ level. Thus, the sum $\sum_{l=1}^{L}h_{l}(\phi)$ “corrects” for the difference between $\ell_{\text{MC}}(\phi)$ (which is based on high-fidelity simulations) and $h_{0}(\phi)$ (which is based on low-fidelity simulations). Note that $f_{\phi}^{l}$ and $f_{\phi}^{l-1}$ share the same random samples $u_{1:m}^{l}$, i.e., they are seed matched. This ensures that $f_{\phi}^{l}$ and $f_{\phi}^{l-1}$ are coupled and highly correlated, leading to a reduction in the variance of $h_{l}(\phi)$ term (Owen, 2013, Ch. 8).
@@
There are several straightforward extensions of our approach which are not covered above so as to not overload notation, but which we now briefly mention. Firstly, each simulator could have its own base measure $\mathbb{U}^{l}$, which could be defined on spaces $\{\mathcal{U}^{l}\}_{l=1}^{L}$ of different dimensions. This is not a problem since we could simply consider $\mathbb{U}$ to be the tensor product measure and $\mathcal{U}$ to be the corresponding tensor product space, in which case all equations above remain valid. However, to ensure the best possible performance, it will still be essential to seed match random numbers where there is overlap; see Owen (2013, Ch. 8) for more details on the use of common random numbers.
@@
Optimisation of the MLMC objective.
Optimising the MLMC loss $\ell_{\text{MLMC}}(\phi)$ in 2 can be challenging due to the conflicting terms $\{\sum_{i=1}^{n_{l}}f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l}),\sum_{i=1}^{n_{l+1}}-f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})\}_{l=0}^{L-1}$. These “positive” and “negative” terms approximate the same expectation $\mathbb{E}[f^{l}_{\phi}]$, but with opposite signs and different number of samples. Later on in the optimisation, as their gradients shrink, one of the terms typically begins to dominate, creating opposing gradient directions. This dynamic leads to unstable updates and, ultimately, causes divergence. To mitigate this issue, we aim for $\nabla_{\phi}h_{0}(\phi)$ to provide a coarse approximation of $\nabla_{\phi}\ell(\phi)$, with the remaining levels serving as corrections. However, under naive optimisation, this breaks down due to imbalanced evolution of gradient throughout the training. To restore the balance, we first rescale the components of $\{\nabla_{\phi}h_{l}(\phi)\}_{l=1}^{L}$ to enforce their corrective role, by making sure that the two opposing components within each $\nabla_{\phi}h_{l}(\phi)$ have comparable norms so that their difference remains small and stable. Then, we apply the gradient surgery technique of Liu et al. (2020), projecting the gradients of $h_{0}(\phi)$ and $\sum_{l=1}^{L}h_{l}(\phi)$ onto each other’s normal planes to further facilitate the optimisation. A detailed explanation is provided in Appendix A.

### v2

@@ 1 Introduction
Nevertheless, the initial training phase of neural SBI methods typically still requires a large number of simulations, preventing their application on computationally expensive (and often more realistic) models which can take hours of compute time for simulating a single data-point. Examples include most tsunami (Behrens and Dias, 2015), wind farm (Kirby et al., 2023), nuclear fusion (Hoppe et al., 2021) and cosmology (Jeffrey et al., 2025) simulators. One avenue to mitigating this issue is multi-fidelity methods (Peherstorfer et al., 2018): we often have access to a sequence of simulators with increasing computational cost and accuracy which we may be able to use to refine existing methods based on a single simulator.
Figure 1: Low- and high-fidelity cosmological simulations from the CAMELS data (Villaescusa-Navarro et al., 2023) studied in Section˜5.4.
This setting is quite common. One example is simulators requiring the numerical solution of ordinary, partial, or stochastic differential equations, where the choice of mesh size or step-size affects both the accuracy of the solution and the computational cost. Another example arises when modelling complex physical, chemical, or biological processes, where low-fidelity simulators can be obtained by neglecting certain aspects of the system. This is exemplified in Figure˜1, where a high-fidelity cosmological simulation includes baryonic astrophysics and thus appear smoother than the low-fidelity simulation.
@@
The key idea behind multi-fidelity methods is to leverage cheaper, less accurate simulations to supplement the more expensive, high-fidelity simulations, ultimately improving efficiency without sacrificing accuracy. This has been popular in the emulation literature since the seminal work of Kennedy and O’Hagan (2000), but applications to SBI are much more recent and include Jasra et al. (2019); Warne et al. (2018); Prescott and Baker (2020, 2021); Warne et al. (2022); Prescott et al. (2024), who proposed multi-fidelity versions of ABC. More recently, (Krouglova et al., 2025) also proposed a multi-fidelity method to enhance neural SBI approximations of the posterior based on transfer learning. However, their method is not supported by any theoretical guarantees, and our numerical experiments in Section˜5.2 will demonstrate that it is sensitive to choices of hyperparameters and Krouglova et al. (2025) unfortunately do not provide any guidance for selecting these.
@@
In this paper, we propose a novel multi-fidelity method which is broadly applicable to neural SBI methods. Taking neural likelihood and neural posterior estimation as the main case-studies, we show that our approach is able to significantly reduce the computational cost of the initial training through the use of multilevel Monte Carlo (Giles, 2015; Jasra et al., 2020) estimates of the training objective. The approach has strong theoretical guarantees; our main result (Theorem˜1) directly links the reduction in computational cost to the accuracy of the low fidelity simulators, and we demonstrate (in Theorem˜2) how to best balance the number of simulations at each fidelity level in the process. Importantly, and unlike Krouglova et al. (2025), our method has no additional hyperparameter which directly relate to fidelity. Our extensive experiments on models from finance, synthetic biology, and cosmology also demonstrate the significant computational advantages provided by our method.
@@ 2 Background
We first recall the basics of SBI methods and the related works on reducing computational cost in Section˜2.1, then provide a brief introduction to multilevel Monte Carlo in Section˜2.2.
@@ 2.1 Simulation-based inference
Let $\{\mathbb{P}_{\theta}\}_{\theta\in\Theta}$ be a parametric family of distributions on some space $\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$ parameterised by $\theta\in\Theta\subseteq\mathbb{R}^{d_{\Theta}}$. We assume that this model is available in the form of a computer code, i.e. as a simulator, wherein simulating from $\mathbb{P}_{\theta}$ is straightforward, but the likelihood function $p(\cdot\,|\,\theta)$ associated with $\mathbb{P}_{\theta}$ is intractable. Simulators can be characterised by a pair $(\mathbb{U},G_{\theta})$, where $\mathbb{U}$ is a distribution (typically simple, such as a uniform or a Gaussian) on a space $\mathcal{U}\subseteq\mathbb{R}^{d_{\mathcal{U}}}$ which captures all of the randomness, and $G_{\theta}:\mathcal{U}\mapsto\mathcal{X}$ is a (deterministic) parametric map called the generator. Simulating $x\sim\mathbb{P}_{\theta}$ can be achieved by first simulating $u\sim\mathbb{U}$, and then applying the generator $x=G_{\theta}(u)$. In this paper, we consider Bayesian inference for the parameters $\theta$ of this simulator-based model given independent and identically distributed (iid) data $x_{1:m}^{o}=\{x_{j}^{o}\}_{j=1}^{m}\in\mathcal{X}^{m}$ collected from some data-generating process. Specifically, we are interested in approximating the posterior with density $\pi(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}p(x_{j}^{o}\,|\,\theta)\pi(\theta)$, where $\pi(\theta)$ is the prior density. As introduced below, this can be achieved via a neural SBI method which approximates the likelihood or posterior.
@@
NLEs (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022; Radev et al., 2023a) are extensions of the synthetic likelihood approach (Wood, 2010; Price et al., 2018) that use flexible conditional density estimators, typically normalising flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), as surrogate models for the likelihood function associated with $\mathbb{P}_{\theta}$. The surrogate conditional density $q_{\phi}^{\text{NLE}}:\mathcal{X}\times\Theta\rightarrow[0,\infty)$ where $q_{\phi}^{\text{NLE}}(\cdot\,|\,\theta)$ is a density function for each $\theta\in\Theta$ and $\phi\in\Phi\subseteq\mathbb{R}^{d_{\Phi}}$ denotes its learnable parameters, is trained by minimising the negative log-likelihood with respect to $\phi$ on simulated samples. More precisely, let $\{(\theta_{i},x_{i})\}_{i=1}^{n}$ be the training data such that $\theta_{i}\sim\pi$ are realisations from the prior and $x_{i}\sim\mathbb{P}_{\theta_{i}}$ are realisations from the simulator, then $\hat{\phi}_{\text{MC}}:=\operatorname*{arg\,min}_{\phi\in\Phi}\ell_{\text{MC}}^{\text{NLE}}(\phi)$, where $\ell_{\text{MC}}^{\text{NLE}}$ is an empirical (Monte Carlo) estimate of negative expected log-density:
@@
Once the surrogate likelihood is trained, Markov chain Monte Carlo (MCMC) or variational inference methods are used to sample from the (approximate) posterior distribution $\pi_{\text{NLE}}(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}q^{\text{NLE}}_{\hat{\phi}_{\text{MC}}}(x^{o}_{j}\,|\,\theta)\pi(\theta)$. NLEs can therefore be regarded as being partially amortised—the surrogate $q_{\phi}^{\text{NLE}}$ need not be trained for a new observed dataset, however, MCMC needs to be carried out again to obtain the new posterior.
For a computationally costly simulator, we note that obtaining training samples can become a bottleneck, which affects the accuracy of estimating the expected loss. Thus, estimating the loss accurately with fewer samples is key to handling costly simulators.
@@
Instead of learning a surrogate likelihood, NPEs learn a mapping $x\mapsto p(\theta\,|\,x)$ from the data to the posterior using conditional density estimators. These are often based on Gaussian mixture network (Bishop, 1994; Papamakarios and Murray, 2016) or normalising flows (Dinh et al., 2014; Papamakarios et al., 2017; Radev et al., 2022). Similar to NLE, the conditional density $q_{\phi}^{\text{NPE}}:\Theta\times\mathcal{X}^{m}\rightarrow[0,\infty)$ is trained by minimising the negative log likelihood with respect to $\phi$ using data $\{(\theta_{i},x_{1:m,i})\}_{i=1}^{n}$ generated by first sampling from the prior $\theta_{i}\sim\pi$ and then the simulator $x_{1:m,i}=(x_{1,i},\ldots,x_{m,i})\sim\mathbb{P}_{\theta_{i}}$:
@@
Once $\phi$ is estimated, the NPE posterior is obtained as $\pi_{\text{NPE}}(\theta\,|\,x_{1:m}^{o})=q^{\text{NPE}}_{\hat{\phi}_{\text{MC}}}(\theta\,|\,x_{1:m}^{o})$. Although training $q_{\phi}^{\text{NPE}}$ incurs an upfront cost, this is a one-time cost as approximate posteriors for new observed datasets are obtained by a simple forward pass of $x_{1:m}^{o}$ through the trained networks, making NPEs fully amortised (in contrast with the partial amortisation of NLE). Similarly to NLE, the computationally costly step in NPE is the generation of training samples from running expensive simulators. Note that both $q^{\text{NLE}}_{\phi}$ and $q^{\text{NPE}}_{\phi}$ usually include a summary function (often architecturally implicit in NPE). This is helpful when $\mathcal{X}$ is high-dimensional or the number of observations $m$ is large (Alsing et al., 2018; Radev et al., 2022), and for NPE it allows conditioning on datasets of different sizes. Recently, alternative training objectives for NPE based on flow matching (Wildberger et al., 2023), diffusion (Geffner et al., 2023; Sharrock et al., 2024; Gloeckler et al., 2024), and consistency models (Schmitt et al., 2024b) have been proposed.
Related work.
We briefly note that beyond the aforementioned multi-fidelity methods, other works also aim to reduce the computational cost of SBI. In the context of ABC, adaptive sampling of the posterior using either sequential Monte Carlo techniques (Sisson et al., 2007; Beaumont et al., 2009; D. Moral et al., 2011) or Gaussian process surrogates (Gutmann and Corander, 2016; Meeds and Welling, 2014) has been a popular approach. A similar approach has been applied to neural SBI (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Papamakarios et al., 2019; Hermans et al., 2020; Durkan et al., 2020), where sequential training schemes are employed to reduce the number of calls to the simulator. Other works have tackled this problem using cost-aware sampling (Bharti et al., 2025), side-stepping high-dimensional estimation (Jeffrey and Wandelt, 2020), early stopping of simulations (Prangle, 2016), dependent simulations (Niu et al., 2023; Bharti et al., 2023), expert-in-the-loop methods (Bharti et al., 2022b), self-consistency properties (Schmitt et al., 2024a), parallelisation of computations (Kulkarni and Moritz, 2023), and the Markovian structure of certain simulators (Gloeckler et al., 2025). Our proposed method, introduced in section˜3, can be combined with all of these compute-efficient SBI methods and is hence complementary to them.
@@ 3 Methodology
Recall that we are performing inference for a simulator $(G_{\theta},\mathbb{U})$ with a prior $\pi$ on the parameter $\theta\in\Theta$. We reparameterise the NLE and the NPE objective in terms of $u$ instead of $x$ (akin to the reparametrisation trick in variational inference), and express them more broadly using an arbitrary loss $\ell:\Phi\rightarrow\mathbb{R}$ (to represent either $\ell^{\text{NLE}}$ or $\ell^{\text{NPE}}$) and an arbitrary function $f_{\phi}:\mathcal{U}^{m}\times\Theta\rightarrow\mathbb{R}$ (to represent either $f^{\text{NLE}}_{\phi}$ or $f^{\text{NPE}}_{\phi}$) as:
@@
The MC estimator of the loss $\ell(\phi)$ is given by ${\ell}_{\text{MC}}(\phi):=\frac{1}{n}\sum_{i=1}^{n}f_{\phi}(u_{1:m,i},\theta_{i})$. Note that the variance of this MC estimator depends on the number of iid samples $n$. Hence, a small $n$ owing to a computationally expensive simulator will lead to a poor estimator for the loss. Now suppose that we have access to a sequence of generators $G^{0}_{\theta}(u),\dots,G^{L-1}_{\theta}(u)$ with varying fidelity levels. Then, for $l=0,\ldots,L-1$, we can define a corresponding sequence of functions $f^{l}_{\phi}:\mathcal{U}^{m}\times\Theta\rightarrow\mathbb{R}$:
$f^{\text{NLE},l}_{\phi}(u_{1:m},\theta):=-\log q^{\text{NLE}}_{\phi}\hskip-2.15277pt\left(G^{l}_{\theta}(u)\,|\,\theta\right),\,\text{and}\,f^{\text{NPE},l}_{\phi}(u_{1:m},\theta):=-\log q^{\text{NPE}}_{\phi}\hskip-2.15277pt\left(\theta\,|\,G^{l}_{\theta}(u_{1}),\ldots,G^{l}_{\theta}(u_{m})\right),$
@@
Equation ˜1 follows by adding then subtracting some terms, and using linearity of expectations. Now suppose that we can simulate from each of these approximate simulators to obtain:
@@
This corresponds to an MLMC estimator of our original objective in ˜1. The first term, $h_{0}(\phi)$, approximates $\ell(\phi)$, but is biased since it uses $f^{0}_{\phi}$ (i.e. the lowest fidelity simulator) rather than $f^{L}_{\phi}$ (the highest fidelity simulator). The additional terms, $h_{1}(\phi),\ldots,h_{L}(\phi)$, correct this bias by estimating the expected difference between the objectives at consecutive fidelity levels.
Within each term $h_{l}(\phi)$, the functions $f_{\phi}^{l}$ and $f_{\phi}^{l-1}$ are evaluated on the same samples $u_{1:m}^{l}$ and $\theta^{l}$, i.e., they are seed-matched. This ensures that the terms are coupled and highly correlated, which leads to a reduction in variance (Owen, 2013, Ch. 8) since
$\text{Var}[h_{l}(\phi)]$ (3)
$=\frac{1}{n_{l}}\left(\text{Var}[f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})]+\text{Var}[f^{l-1}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})]-2\text{Cov}\left[f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l}),~f^{l-1}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})\right]\right),$
and the covariance will be large. This covariance will be particularly large the more similar the two functions $f^{l}_{\phi}$ and $f^{l-1}_{\phi}$ are, and we therefore expect $\text{Var}[h_{l}(\phi)]$ to be smallest in those settings. We can also immediately see that without seed-matching, the covariance will be small and the variance will be large, highlighting why seed-matching is essential for MLMC.
@@
There are several straightforward extensions of our approach which are not covered above so as to not overload notation. Firstly, each simulator could have its own base measure $\mathbb{U}^{l}$, which could be defined on spaces $\{\mathcal{U}^{l}\}_{l=1}^{L}$ of different dimensions. This is not a problem since we could simply consider $\mathbb{U}$ to be the tensor product measure and $\mathcal{U}$ to be the corresponding tensor product space, in which case all equations above remain valid. Similarly, the parameter space may differ across simulators. However, to ensure the best possible performance, it will still be essential to seed-match random numbers where there is overlap; see Owen (2013, Ch. 8) for more details on the use of common random numbers, and Section˜5 for a study of this issue for multilevel neural SBI.
@@
Optimisation.
Algorithm 1 MLMC gradient adjustment
Input $\nabla_{\phi}h_{0}(\phi),\{\zeta_{\phi}^{l,+},\zeta_{\phi}^{l-1,-}\}_{l=1}^{L}$, $\epsilon>0\approx 0$
# Rescaling the gradients
for $l=1$ to $L$
[…]
end for
$\nabla_{\phi}h_{c}(\phi)\leftarrow\sum_{l=1}^{L}(\zeta_{\phi}^{l,+}+\zeta_{\phi}^{l-1,-})$
# Projecting the gradients (only when conflicting)
if $\nabla_{\phi}h_{0}(\phi)\cdot\nabla_{\phi}h_{c}(\phi)<0$ then
[…]
else $\tilde{\nabla}_{\phi}h_{0}(\phi)\leftarrow\nabla_{\phi}h_{0}(\phi),\tilde{\nabla}_{\phi}h_{c}(\phi)\leftarrow\nabla_{\phi}h_{c}(\phi)$
end if
Output $\tilde{\nabla}_{\phi}\ell_{\text{MLMC}}(\phi)\leftarrow\tilde{\nabla}_{\phi}h_{0}(\phi)+\tilde{\nabla}_{\phi}h_{c}(\phi)$
Gradient-based optimisation of the MLMC objective $\ell_{\text{MLMC}}(\phi)$ can be challenging due to the ‘conflicting’ gradients which appear in consecutive terms $\nabla_{\phi}h_{l}(\phi)$ and $\nabla_{\phi}h_{l+1}(\phi)$. More precisely, the term $\nabla_{\phi}h_{l}(\phi)$ always contains
$\zeta^{l,+}_{\phi}:=\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}\nabla_{\phi}f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l}),$
which approximates $\nabla_{\phi}\mathbb{E}[f^{l}_{\phi}]$, whilst the term $\nabla_{\phi}h_{l+1}(\phi)$ always contains
$\zeta_{\phi}^{l,-}:=-\frac{1}{n_{l+1}}\sum_{i=1}^{n_{l+1}}\nabla_{\phi}f^{l}_{\phi}(u_{1:m,i}^{l+1},\theta_{i}^{l+1}),$
which approximates $\nabla_{\phi}\mathbb{E}[f^{l}_{\phi}]$. In the infinite-sample limit, $\nabla_{\phi}\mathbb{E}[f^{l}_{\phi}]$ and $\nabla_{\phi}\mathbb{E}[f^{l}_{\phi}]$ cancel out, but this is not the case for $\zeta_{\phi}^{l,+}$ and $\zeta_{\phi}^{l,-}$ since we are typically working with only a small number of expensive simulations. When naively applying standard gradient-based optimisation methods on this loss, we observe that the training dynamics is typically dominated by only one of these two quantities until approaching stationarity, at which point the conflicting gradients lead to unstable updates and, ultimately, often cause divergence.
To mitigate this issue, we use a combination of gradient adjustments summarised in Algorithm 1. Firstly, we rescale $\zeta_{\phi}^{l,+}$ and $\zeta_{\phi}^{l,-}$ to ensure that they have comparable norms and that their difference remains small and stable. Secondly, we apply the gradient projection technique of Liu et al. (2020), projecting the gradients of $h_{0}(\phi)$ and $h_{c}(\phi):=\sum_{l=1}^{L}h_{l}(\phi)$ onto each other’s normal planes to reduce the impact of conflicting gradients. We observe empirically that combining these two techniques significantly improves the stability of the optimisation throughout the training and leads to better performance; see Section˜B.6 for a detailed comparison.